# L-equivalence and Fourier–Mukai partners of cubic fourfolds

### v1

@@ 5. Implications of $\mathbb{L}$-equivalence
The following proposition is the analogue for cubic fourfolds of [Efi18, Proposition 3.3]. We remark that under Assumption 3.1 this follows immediately from [Huy16, Theorem 1.1].
@@
Denote by $S_{D}$ the set of all symmetric pairings $T(H)\otimes T(H)\to\mathbb{Z}(-2)$ which are morphisms in $\operatorname{HS}_{\mathbb{Z},4}$ and have discriminant $D$. The isomorphism $f(1)$ pulls back the intersection form on $H^{4}(Y,\mathbb{Z})$, and restricting to the transcendental part gives an element of $S_{D}$.
Consider the group $G:=\operatorname{Aut}_{\operatorname{HS}_{\mathbb{Z},2}}(T(X))$ with its natural action on $S_{D}$. Up to Hodge automorphisms of $T(X)$, the possible transcendental intersection forms coming from such $Y$ are parametrized by the $G$-orbits in $S_{D}$.
The finiteness of the $G$-orbits in $S_{D}$ now follows exactly as in [Efi18, Proposition 3.3]. For each of these finitely many orbits we can apply Lemma 5.2 and get finitely many isomorphism classes of cubic fourfold $Y$, concluding the proof. ∎
@@
Assume first we have a Hodge isometry $\phi\colon H^{4}(X,\mathbb{Z})\to H^{4}(Y,\mathbb{Z})$, and consider the class $v\coloneqq\phi^{-1}(h_{Y}^{2})$. Since $\phi$ is a Hodge morphism and $h_{Y}^{2}\in A(Y)$, we have $v\in A(X)$, and since $\phi$ is an isometry $(v,v)=(h_{Y}^{2},h_{Y}^{2})=3$. The lattice $A(X)$ is positive definite. In a positive definite integral lattice, there are only finitely many vectors of a given square, see [Hanke2013, Section 2.1] for a reference. Hence the set $S_{3}:=\{v\in A(X):(v,v)=3\}$ is finite. Fix $v\in S_{3}$ and assume $Y_{1}$ and $Y_{2}$ are cubic fourfolds with Hodge isometries $\phi_{i}\colon H^{4}(X,\mathbb{Z})\to H^{4}(Y_{i},\mathbb{Z})$ such that $\phi_{i}^{-1}(h_{Y_{i}}^{2})=v$. The composition $\phi_{2}\circ\phi_{1}^{-1}$ is a Hodge isometry between $H^{4}(Y_{1},\mathbb{Z})$ and $H^{4}(Y_{2},\mathbb{Z})$. Moreover, this composition sends $h_{1}^{2}$ to $h_{2}^{2}$ and so gives an identification $H^{4}(Y_{1},\mathbb{Z})_{\operatorname{prim}}\cong H^{4}(Y_{2},\mathbb{Z})_{\operatorname{prim}}$. By Theorem 2.9 we have $Y_{1}\cong Y_{2}$.
@@
The following theorem is the analogue for cubic fourfolds of [Efi18, Theorem 3.4]. Again, under Assumption 3.1, Theorem 5.4 implies that $\mathbb{L}$-equivalent cubic fourfolds are Fourier-Mukai partners, and the results again follows from [Huy16, Theorem 1.1].
@@
Let $Y$ be a cubic fourfold such that $X\sim_{\mathbb{L}}Y$, and consider the integral Hodge realisation $\operatorname{Hdg}_{\mathbb{Z}}$ constructed in [Efi18, Section 3]. From $[X]=[Y]$ in $K_{0}(\operatorname{Var}_{\mathbb{C}})[\mathbb{L}^{-1}]$ we deduce $H^{4}(X,\mathbb{Z})]=[H^{4}(Y,\mathbb{Z})]$ in $K_{0}(\operatorname{HS}^{4}_{\mathbb{Z}})$.
By [Efi18, Theorem 2.7], applied to the additive category $\operatorname{HS}^{4}_{\mathbb{Z}}$, the class $H^{4}(X,\mathbb{Z})]$ in $K_{0}(\operatorname{HS}^{4}_{\mathbb{Z}})$ contains only finitely many isomorphism classes of objects. In particular there are only finitely many possibilities for $H^{4}(Y,\mathbb{Z})$ as an integral Hodge structure when $Y$ ranges over cubic fourfolds with $X\sim_{\mathbb{L}}Y$. The thesis then follows from Proposition 5.1. ∎
The following result appears as Theorem 1.4 in [BB2025].
@@
Let $X$ and $Y$ be cubic fourfolds satisfying Assumption 3.1 (for example, if $X$ is a cubic fourfold satisfying one of the properties of Proposition 3.3) Moreover, assume $\operatorname{rk}(T(X))\neq 4$ and $\operatorname{End}_{\text{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, if $X$ and $Y$ are $\mathbb{L}$-equivalent, they are Fourier–Mukai partners.
@@
The following result appears as Theorem 1.2 in [BB2025].
@@
Let $X$ be a cubic fourfold not contained in any of the Hassett divisors. Assume moreover that $\operatorname{End}_{\text{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, for any cubic fourfold $Y$, we have
@@
The following result appears as Proposition 4.4 in [BB2025].
@@
Since $X$ and $Y$ are assumed to be $\mathbb{L}$-equivalent, $X^{[2]}$ and $Y^{[2]}$ are $\mathbb{L}$-equivalent by [Oka21, Lemma 2.1]. Now the result follows directly from the beautiful formula of Galkin–Shinder [GS14]: for any cubic fourfold $X$, we have
@@
Let $X$ and $Y$ be cubic fourfolds. Assume that $X$ satisfies Assumption 3.1, that $\operatorname{rk}(T(X))\neq 4$, and that $\operatorname{End}_{\text{Hodge}}(T(X))\simeq\mathbb{Z}$. Then if $F(X)\sim_{\mathbb{L}}F(Y)$, there is an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$, hence also an equivalence $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$.
@@
By [Mei25, Theorem 3.5] and Lemma 2.19, the gluing group
@@
has order $|G|=\operatorname{disc}(T(F(X)))$. Therefore, if $F(X)$ is $\mathbb{L}$-equivalent to $F(Y)$, it follows from Lemma 2.17 that $\operatorname{disc}(T(F(X)))=\operatorname{disc}(T(F(Y)))$. Since $T(F(X))$ is Hodge anti-isometric to $T(X)$, we find that $[T(X)(1)]=[T(Y)(1)]$. By Theorem 2.20, there is a Hodge isometry $T(X)\simeq T(Y)$, hence an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ by Assumption 3.1. The fact that there is an equivalence $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$ follows from the main result of [KS25]. ∎
@@
By [Huy1999, Theorem 4.6] $X$ and $Y$ are special fibres of two one-dimensional families isomorphic over a punctured disk. Near the central fibre the generic fibres are isomorphic, hence define the same class in $K_{0}(\operatorname{Var}_{\mathbb{C}})$. Then the result follows from the motivic measure $\operatorname{Vol}_{\mathbb{C}}$ described in [NS2019, Theorem 3.1.1]. ∎
@@
Let $X$ and $Y$ be cubic fourfolds. If $F(X)$ is birational to $F(Y)$, then $F(X)$ and $F(Y)$ are $L$-equivalent. In other words, we have BF $\implies$ LF.

### v2

@@ 5. Implications of $\mathbb{L}$-equivalence
The following proposition is the analogue for cubic fourfolds of [11, Proposition 3.3]. We remark that under Assumption 3.1 this follows immediately from [21, Theorem 1.1].
@@
Denote by $S_{D}$ the set of all symmetric pairings $T(H_{X})\otimes T(H_{X})\to\mathbb{Z}(-2)$ which are morphisms in $\operatorname{HS}_{\mathbb{Z},4}$ and have discriminant $D$. The isomorphism $f(1)$ pulls back the intersection form on $H^{4}(Y,\mathbb{Z})$, and restricting to the transcendental part gives an element of $S_{D}$.
Consider the endomorphism ring $R:=\operatorname{End}_{\operatorname{HS}_{\mathbb{Z},2}}(T(H_{X}))$, and the group $G:=\operatorname{Aut}_{\operatorname{HS}_{\mathbb{Z},2}}(T(H_{X}))$ with its natural action on $S_{D}$. Up to Hodge automorphisms of $T(H_{X})$, the possible transcendental intersection forms coming from such $Y$ are parametrized by the $G$-orbits in $S_{D}$.
Now we need to show that the number of $G$-orbits in $S_{D}$ are finite. We follow the same proof as in [11, Proposition 3.3]. The only change is that we use $T(H_{X})$ instead of the transcendental lattice of a K3 surface. For carrying on the proof we need to check that $T(H_{X})$ satisfies the hypotheses of Zarhin’s theorem [54] (namely that $T(H_{X})$ is a polarized weight $2$ Hodge structure of K3 type). As a consequence, the $\mathbb{Q}$-algebra $R\otimes\mathbb{Q}$ is a totally real number field, or an imaginary quadratic extension of a totally real number field. Denote by $\beta$ an element of $S_{D}$ and consider the associated Hodge endomorphism $a(\beta)\in F$. Efimov’s argument then applies verbatim and gives the finiteness of $S_{D}/G$.
Finally, for each of these finitely many orbits we can apply Lemma 5.2 and get finitely many isomorphism classes of cubic fourfold $Y$, concluding the proof. ∎
@@
Assume first we have a Hodge isometry $\phi\colon H^{4}(X,\mathbb{Z})\to H^{4}(Y,\mathbb{Z})$, and consider the class $v\coloneqq\phi^{-1}(h_{Y}^{2})$. Since $\phi$ is a Hodge morphism and $h_{Y}^{2}\in A(Y)$, we have $v\in A(X)$, and since $\phi$ is an isometry $(v,v)=(h_{Y}^{2},h_{Y}^{2})=3$. The lattice $A(X)$ is positive definite. In a positive definite integral lattice, there are only finitely many vectors of a given square, see [16, Section 2.1] for a reference. Hence the set $S_{3}:=\{v\in A(X):(v,v)=3\}$ is finite. Fix $v\in S_{3}$ and assume $Y_{1}$ and $Y_{2}$ are cubic fourfolds with Hodge isometries $\phi_{i}\colon H^{4}(X,\mathbb{Z})\to H^{4}(Y_{i},\mathbb{Z})$ such that $\phi_{i}^{-1}(h_{Y_{i}}^{2})=v$. The composition $\phi_{2}\circ\phi_{1}^{-1}$ is a Hodge isometry between $H^{4}(Y_{1},\mathbb{Z})$ and $H^{4}(Y_{2},\mathbb{Z})$. Moreover, this composition sends $h_{1}^{2}$ to $h_{2}^{2}$ and so gives an identification $H^{4}(Y_{1},\mathbb{Z})_{\operatorname{prim}}\cong H^{4}(Y_{2},\mathbb{Z})_{\operatorname{prim}}$. By Theorem 2.9 we have $Y_{1}\cong Y_{2}$.
@@
The following theorem is the analogue for cubic fourfolds of [11, Theorem 3.4]. Again, under Assumption 3.1, Theorem 5.4 implies that $\mathbb{L}$-equivalent cubic fourfolds are Fourier-Mukai partners, and the results again follows from [21, Theorem 1.1].
@@
Let $Y$ be a cubic fourfold such that $X\sim_{\mathbb{L}}Y$, and consider the integral Hodge realisation $\operatorname{Hdg}_{\mathbb{Z}}$ constructed in [11, Section 3]. From $[X]=[Y]$ in $K_{0}(\operatorname{Var}_{\mathbb{C}})[\mathbb{L}^{-1}]$ we deduce $H^{4}(X,\mathbb{Z})]=[H^{4}(Y,\mathbb{Z})]$ in $K_{0}(\operatorname{HS}^{4}_{\mathbb{Z}})$.
By [11, Theorem 2.7], applied to the additive category $\operatorname{HS}^{4}_{\mathbb{Z}}$, the class $H^{4}(X,\mathbb{Z})]$ in $K_{0}(\operatorname{HS}^{4}_{\mathbb{Z}})$ contains only finitely many isomorphism classes of objects. In particular there are only finitely many possibilities for $H^{4}(Y,\mathbb{Z})$ as an integral Hodge structure when $Y$ ranges over cubic fourfolds with $X\sim_{\mathbb{L}}Y$. The thesis then follows from Proposition 5.1. ∎
The following result appears as Theorem 1.4 in [4].
@@
Let $X$ and $Y$ be cubic fourfolds satisfying Assumption 3.1 (for example, if $X$ is a cubic fourfold satisfying one of the properties of Proposition 3.3) Moreover, assume $\operatorname{rk}(T(X))\neq 4$ and $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, if $X$ and $Y$ are $\mathbb{L}$-equivalent, they are Fourier–Mukai partners.
@@
The following result appears as Theorem 1.2 in [4].
@@
Let $X$ be a cubic fourfold not contained in any of the Hassett divisors. Assume moreover that $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, for any cubic fourfold $Y$, we have
@@
The following result appears as Proposition 4.4 in [4].
@@
Since $X$ and $Y$ are assumed to be $\mathbb{L}$-equivalent, $X^{[2]}$ and $Y^{[2]}$ are $\mathbb{L}$-equivalent by [46, Lemma 2.1]. Now the result follows directly from the beautiful formula of Galkin–Shinder [14]: for any cubic fourfold $X$, we have
@@
Let $X$ and $Y$ be cubic fourfolds. Assume that $X$ satisfies Assumption 3.1, that $\operatorname{rk}(T(X))\neq 4$, and that $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then if $F(X)\sim_{\mathbb{L}}F(Y)$, there is an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$, hence also an equivalence $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$.
@@
By [40, Theorem 3.5] and Lemma 2.19, the gluing group
@@
has order $|G|=\operatorname{disc}(T(F(X)))$. Therefore, if $F(X)$ is $\mathbb{L}$-equivalent to $F(Y)$, it follows from Lemma 2.17 that $\operatorname{disc}(T(F(X)))=\operatorname{disc}(T(F(Y)))$. Since $T(F(X))$ is Hodge anti-isometric to $T(X)$, we find that $[T(X)(1)]=[T(Y)(1)]$. By Theorem 2.20, there is a Hodge isometry $T(X)\simeq T(Y)$, hence an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ by Assumption 3.1. The fact that there is an equivalence $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$ follows from the main result of [29]. ∎
@@
By [22, Theorem 4.6] there exist two smooth projective families $\mathcal{X}$ and $\mathcal{Y}$ over a smooth one dimensional base $S$ with special fibres $\mathcal{X}_{0}\cong X$ and $\mathcal{Y}_{0}\cong Y$, such that $\mathcal{X}$ and $\mathcal{Y}$ are isomorphic over $S\smallsetminus\{0\}$.
Now consider the generic fibres over $S\smallsetminus\{0\}$. As in [43] let $R$ be the completion of the local ring of $S$ at 0 and let $K=\operatorname{Frac}(R)$, so that $\operatorname{Spec}K$ is the generic point of $\operatorname{Spec}(R)$. By taking base changes we get the restrictions $\mathcal{X}_{R}$ and $\mathcal{Y}_{R}$ of $\mathcal{X}$ and $\mathcal{Y}$ to a formal neighbourhood of 0, and the corresponding generic fibres $\mathcal{X}_{K}$ and $\mathcal{Y}_{K}$.
Recall that we have a canonical identification of special fibres $(\mathcal{X}_{R})_{k}=\mathcal{X}_{R}\times_{\operatorname{Spec}R}\operatorname{Spec}k\cong\mathcal{X}\times_{S}\operatorname{Spec}k\cong\mathcal{X}_{0}$, and similarly $(\mathcal{Y}_{R})_{k}\cong\mathcal{Y}_{0}$.
The isomorphism of the families over $S\smallsetminus\{0\}$ induces an isomorphism over $K$ of $\mathcal{X}_{K}$ and $\mathcal{Y}_{K}$, which induces $\mathcal{X}_{K}]=[\mathcal{Y}_{K}]$ in $K_{0}(\operatorname{Var}_{K})$.
Now apply the motivic volume morphism $\operatorname{Vol}_{K}:K_{0}(\operatorname{Var}_{K})\to K_{0}^{\hat{\mu}}(\operatorname{Var}_{\mathbb{C}})$ of [43, Theorem 3.1.1]. From $\mathcal{X}_{K}]=[\mathcal{Y}_{K}]$ we get $\operatorname{Vol}_{K}([\mathcal{X}_{K}])=\operatorname{Vol}_{K}([\mathcal{Y}_{K}])$. From the fact that $\mathcal{X}_{R}$ ($\mathcal{Y}_{R}$) is a smooth proper $R$ model of $\mathcal{X}_{K}$ ($\mathcal{Y}_{K}$), so $\operatorname{Vol}_{K}([\mathcal{X}_{K}])=[\mathcal{X}_{0}]=[X]$ and $\operatorname{Vol}_{K}([\mathcal{Y}_{K}])=[\mathcal{Y}_{0}]=[Y]$. This concludes the proof. ∎
@@
Let $X$ and $Y$ be cubic fourfolds. If $F(X)$ is birational to $F(Y)$, then $F(X)$ and $F(Y)$ are $L$-equivalent. In other words, we have $\mathbf{BF}\implies\mathbf{LF}$.